Rita is 23, and joel is 15. How many years ago was rita 3 times as old joel?

Found 2 solutions by ikleyn, greenestamps:
Answer by ikleyn(52794)

.

    23 - x = 3*(15-x)

    23 - x = 45 - 3x

    3x - x = 45 - 23

    2x     = 22

     x     = 22/2 = 11.


ANSWER.  11 years ago.

Solved.


Answer by greenestamps(13200)



You should know how to set up and solve this problem using formal algebra, as shown by the other tutor.

However, you can get some good brain exercise and good practice in problem solving by solving the problem informally, using logical reasoning and simple arithmetic.

The difference in their ages is, and always will be, 8 years.

Given that age difference, when Rita was 3 times as old as Joel, mental arithmetic tells us Rita was 12 and Joel was 4.

Rita is now 23, so that was 23-12 = 11 years ago.

ANSWER: 11 years